# MoIIE: Mixture of Intra- and Inter-Modality Experts for Large Vision Language Models

### v1

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across multi-modal tasks by scaling model size and training data. However, these dense LVLMs incur significant computational costs and motivate the exploration of sparse Mixture of Experts (MoE) architectures. While MoE improve parameter efficiency, effectively applying MoE to simultaneously model modality-specific features and cross-modal associations in LVLMs remains challenging. In this work, we propose to incorporate Mixture of Intra- and Inter-Modality Experts (MoIIE) to LVLMs. For each token, expert routing is guided by its modality, directing tokens to their respective intra-modality experts as well as a shared pool of inter-modality experts, enabling the model to jointly learn rich intra-modal features and cross-modal interactions. We further introduce an effective and straightforward two-stage training strategy, which facilitates the direct activation of both MoE and multi-modal capabilities. Extensive experiments across different data scales and LLM backbone demonstrate the effectiveness, efficiency and generality of our approach. Notably, our MoIIE models with 5.5B and 11.3B activated parameters match or even surpass the performance of existing advanced open-source MoE-LLMs based multi-modal models that involve more activated parameters. The code is available at https://github.com/AlenjandroWang/MoIIE.
@@ 1 Introduction
To this end, we proposes the Mixture of Intra- and Inter-Modality Experts (MoIIE), to simultaneously capture both modality-specific features and cross-modal associations in LVLMs. As illustrated in Figure 1 (b), MoIIE comprises three distinct groups of experts: two intra-modality groups that specialize in language and vision, independently processing text and image tokens; and an inter-modality expert group shared by both modalities that focuses on cross-modal interactions between text and image tokens. Correspondingly, MoIIE learns two dedicated routers respectively for image and text tokens, with each token dynamically routed to the most relevant experts from the corresponding intra-modality group and the inter-modality group. As not all tokens are strongly associated across modalities, such as function words in text or visual regions without descriptive content, the routing mechanism allows for flexible combinations. Some tokens may activate only intra-modality experts or only inter-modality experts, while others may access both expert groups simultaneously. This dynamic routing strategy is consistently applied to both image and text tokens.
Additionally, for multi-modal MoE training, most existing approaches adopt a sparse upcycling strategy Komatsuzaki et al. (2023) that transforms dense LVLMs into sparse models. This strategy generally follows a three-stage training pipeline: multi-modal pre-training of the dense model, fine-tuning the dense model on diverse multitasking datasets with all parameters unfrozen, and model sparsification with only expert modules trained Lin et al. (2024a); Chen et al. (2024b). However, this pipeline is relatively cumbersome and the limited training during the final stage potentially undermines the model’s generalization capabilities. In contrast, we propose that multi-modal fine-tuning and sparsification can be jointly optimized with all parameters updated in a more straightforward and effective two-stage training strategy. In the first stage, visual inputs are aligned with the LLM backbone through multi-modal pre-training datasets. In the second stage, MoIIE is integrated and the entire model is optimized using fine-tuning datasets, to simultaneously learn (1) versatile multi-modal understanding and instruction-following capabilities, while enabling (2) expert learning of modality-specific knowledge with cross-modal association within the MoIIE module. Our experimental results demonstrate that this simplified strategy is easy to implement and also effectively enhances performance across diverse downstream tasks.
@@ 2.2.2 Sparse MoIIE Forward
$\text{MoIIE}(x_{n})=\begin{cases}\sum_{i=1}^{K}G_{i}^{\mathcal{T}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{T}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{T}\\ \sum_{i=1}^{K}G_{i}^{\mathcal{I}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{I}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{I}\end{cases}$ (2)
where each token $x_{n}$ is processed by its top-$K$ activated experts $\text{E}_{i}(x_{n})$, with their outputs aggregated through a probability-weighted summation. Specifically, image tokens are routed by the visual router $G^{\mathcal{I}}(x)$ to the top-$K$ experts selected from the combined pool of intra-vision and inter-modality groups $E^{\mathcal{I/S}}$, while textual tokens are routed by the textual router $G^{\mathcal{T}}(x)$ to the top-$K$ experts from the intra-language and inter-modality groups $E^{\mathcal{T/S}}$.
The routing mechanism enables flexible expert selection patterns: each token may activate only intra-modality experts, only inter-modality experts, or a combination of both. This hierarchical routing strategy ensures efficient multi-modal learning while simultaneously modeling both modality-specific features and cross-modal associations. All experts in $\mathcal{E}$ are initialized from the LLM’s FFN layers, preserving prior knowledge while enabling adaptive multi-modal specialization.
@@ 3.1 Implementation Details
We employ a pre-trained SigLIP model Zhai et al. (2023b) as the visual encoder and a two-layer MLP as the connection module. For the LLM backbone, we utilize phi-3-mini Marah Abdin (2024) and LLaMA3-8b AI@Meta (2024). Our method and all compared MoE architectures use a 4-experts configuration. Specifically, our MoIIE module includes 2 intra-modality experts (one for vision and one for language) and 2 inter-modality experts. More detailed discussion of expert settings is provided in the ablation studies 3.6.
@@
Vanilla MoE 2M Phi-3-mini 75.1 62.8 81.5 42.8 70.6 86.3 28.8 65.1 42.2 53.1 65.2 40.6 29.9 57.2
Modality MoE 2M Phi-3-mini 75.1 62.9 81.7 41.0 71.0 86.2 29.3 66.0 42.7 53.1 65.5 40.4 31.0 57.4
MoIIE 2M Phi-3-mini 75.3 63.2 81.8 42.8 71.2 86.6 30.5 65.4 42.9 53.3 65.9 41.3 31.9 57.9
With More Visual Instruction Data
@@
MoIIE 2M LLaMA3-8B 75.7 64.9 82.3 47.5 73.0 87.0 36.5 67.9 44.7 56.0 76.9 42.8 32.2 60.6
During training, we utilize the AdamW optimizer Kingma & Ba (2017) with a cosine learning rate scheduler for one epoch across both stages. In the first stage, only the connection module is optimized with a learning rate of $1\times 10^{-3}$, using the Bunny-pretrain-LAION-2M dataset He et al. (2024). In the second stage, all parameters are unfrozen for joint SFT and sparse upcycling training. The learning rate is set to $2\times 10^{-6}$ for the visual encoder and $2\times 10^{-5}$ for all other components. For this stage, we attempt training with different scales of visual instruction data: (1) 1.3M samples from MGM-Instruction Li et al. (2024c); (2) 2M samples by adding Bunny-695K He et al. (2024); (3) 2.7M samples by further incorporating LLaVA-NEXT-779k Liu et al. (2024b). Training is conducted using DeepSpeed Shuaiwen Leon Song (2023) with ZeRO-3 optimization. All compared baselines in the main results and our method follow the same training setup and data configurations. Further implementation details are provided in Appendix A.
@@ 3.2 Evaluation Benchmarks
We conduct a comprehensive evaluation across 13 diverse multi-modal benchmarks. Specifically, general multi-modal benchmarks include MMBench-EN Liu et al. (2024c), MM-Vet Yu et al. (2024), GQA Hudson & Manning (2019), VQAv2 and SEED-Image  Li et al. (2023a). For knowledge-based question answering, we utilize MMMU validation subset Yue et al. (2024), AI2D Kembhavi et al. (2016),SciQA-IMG Lu et al. (2022) and Mathvista Lu et al. (2024b). For OCR-based question answering, we assess performance on TextVQA Singh et al. (2019), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021). Additionally, we evaluate hallucination robustness using POPE Li et al. (2023c) and HallusionBench Guan et al. (2024).
@@ 3.3 Main Results
We compare our MoIIE-equipped model against three baseline variants: the original dense model (Dense) without MoE, the vanilla MoE model MistralAITeam (2023) that routes tokens from all modalities to a shared pool of experts (Vanilla MoE), and the model with modality-specific experts where text and image tokens are routed to their respective specialized experts (Modality MoE). The comparisons are conducted across multiple visual instruction tuning data scales and varying LLM backbones, as shown in Table 1. The key findings are as follows:
(1) The MoIIE module consistently outperforms all other architectural variants across most benchmarks with different data scales and LLM backbone sizes, demonstrating that integrating both modality-specific features and cross-modal associations can effectively enhance performance on diverse multi-modal tasks. Further analysis of expert modules is provided in Appendix A.
@@
(3) We further provide detailed performance trends with increasing fine-tuning data in Figure 3. As the data scale continues to increase, our MoIIE progressively improves while other architectures, especially Dense and Modality MoE, encounter performance limitations. Moreover, with larger training datasets, our MoIIE consistently outperforms all alternatives, with the performance gap widening between MoIIE and both Dense and Modality MoE. This suggests that the MoIIE framework offers superior scaling properties for multi-modal learning, effectively leveraging larger datasets to enhance representation power without the parameter inefficiency of dense models or the limited cross-modal reasoning of strictly modality-separated experts.
@@ 3.4 Comparison with MoE-based LLMs Initialized LVLMs
We further compare our MoIIE pre-trained on the LLaVA-OV-Single-Image dataset Li et al. (2024a) with existing open-sourced LVLMs initialized from MoE-based LLMs to demonstrate its effectiveness. While SPHINX-X Liu et al. (2024a), MGM Li et al. (2024c) and CuMO Li et al. (2024b) that are built on Mixtral-8x7B MistralAITeam (2023) leverage large-scale language data to provide robust initialization and enhanced language capabilities, they typically demand significant computational resources, rendering them less practical for resource-constrained settings. In contrast, our approach achieves comparable performance on diverse multi-modal benchmarks with 17% fewer activated parameters, reducing overall parameter count and lowering deployment and training costs. Furthermore, our approach allows for the construction of MoE-based LVLMs from any LLM backbones, supporting more flexible expert settings without sacrificing performance.
@@
MoE-based LVLMs model with pre-trained MoE LLM backbone
@@
MoE-based LVLMs model from dense LLM backbone
MoIIE-A5B 5.5B 7B 69.5 63.4 82.5 87.1 77.8 42.6 43.6 46.4 90.1
MoIIE-A11B 11.3B 16B 70.5 64.5 82.9 87.3 77.2 47.3 42.7 46.0 89.4
@@ 3.5 The Discussion of Training Pipeline
Converting dense LLMs into MoE-based LVLMs typically involves a three-stage training process Lin et al. (2024a); Chen et al. (2024b). First, multi-modal pre-training that transforms dense LLMs into dense LVLMs by aligning visual and textual representations. The second stage is to fine-tune the dense LVLMs with multi-modal instruction data followed by a multi-modal sparse upcycling phase to obtain MoE-based LVLMs. However, this three-stage training pipeline is relatively cumbersome and the limited training during the final stage potentially undermines the model’s generalization capabilities. In contrast, we propose a streamlined two-stage training strategy that consolidates SFT and model sparsification into a single stage after pretraining, enabling the direct activation of both multi-modal and MoE capabilities.
The three stages training variants utilize 2M visual instruction data for SFT with all parameters tuned, and LLaVA-NEXT-779k for sparse upcycling with only MoE layers tuned. The two stages training variants utilize 2.7M visual instruction data for joint SFT and sparse upcycling to tune all model parameters. As shown in Table 3, our proposed two-stage training method consistently outperforms the previous three-stage training variants under the same training settings across comprehensive multi-modal benchmarks, demonstrating that our approach is simpler and also more effective.
@@ 3.6 Ablation Study
Increasing the number of experts typically improves MoE performance  Lepikhin et al. (2020); FedusF et al. (2022). In Table 4, we vary the total number of experts in the MoIIE module while keeping the number of activated parameters and the allocation ratio of different expert groups constant. Results show that expanding from 4 to 8 experts results in performance gains of 0.3, 0.2, and 0.9 points on MMBench, GQA, MM-Vet, and MMMU, respectively, while performance on POPE and TextVQA remains stable or slightly decreases. These results suggest that a larger expert pool can improve generalization on complex benchmarks, but may require more extensive data to fully leverage its capacity to achieve performance improvement.
@@
Maintaining a balanced allocation of intra-modality and inter-modality experts proves beneficial for performance. As shown in Table 4, we examine the impact of expert balance under an 8-expert configuration. In the Balanced setting, we maintain a 1:1 ratio between intra-modality and inter-modality experts, specifically consisting of 2 vision experts, 2 language experts, and 4 shared experts. Each router selects the top-2 experts from its corresponding intra-modality pool and the inter-modality expert pool, resulting in 2 + 4 candidate experts. In contrast, the unbalanced setting adopts a 3:1 ratio, with 3 vision experts, 3 language experts, and only 2 shared experts. In this configuration, each router selects the top-2 experts from 3 intra-modality and 2 inter-modality candidates. Experimental results indicate that the balanced setting consistently outperforms the unbalanced configuration, suggesting that maintaining an even distribution of intra- and inter- modality experts yields more effective cross-modal interaction and overall performance.
@@
We explore two architectural configurations for integrating the MoIIE module into LVLMs: the “interleaved” configuration and the “full” configuration following the designs in Lin et al. (2024a); AI@Meta (2025). In the interleaved setup, MoIIE module are incorporated every other layer, interspersed with standard dense layers throughout the model, maintaining a balance between sparse and dense computations. In contrast, the full configuration replaces all layers with MoIIE layers, maximizing the use of sparse experts across the entire network.
As shown in Table 4, fully replacing all layers with MoIIE modules does not necessarily lead to significant performance improvements but substantially increases training costs and computational demands due to the larger parameter count. In contrast, the interleaved configuration achieves comparable performance while maintaining higher training efficiency and reduced resource consumption. Therefore, we adopt the interleaved configuration as the optimal design choice for MoIIE-based LVLMs, ensuring a well-balanced trade-off between performance, efficiency, and scalability.
The Impact of Load Balance Coefficient
In MoE architectures, there is an inherent trade-off between load balance and model performance. The auxiliary loss is designed to promote load balance among experts Cai et al. (2024). However, we observe that while enforcing balance, this loss inevitably interferes with multi-modal learning, potentially hindering the model’s capacity to capture cross-modal associations. As shown in Table 4, we evaluate MoIIE under different values of the auxiliary loss weight $\alpha$. The results indicate that increasing $\alpha$ leads to a noticeable decline in overall performance, particularly on MMMUval and MM-Vet, where the drop is most pronounced. This suggests that finding an optimal balance is essential to maintain robust multi-modal understanding, ensuring that expert specialization is not excessively constrained by strict load balancing.

### v2

@@ Abstract
Large Vision-Language Models (LVLMs) have demonstrated remarkable performance across multi-modal tasks by scaling model size and training data. However, these dense LVLMs incur significant computational costs and motivate the exploration of sparse Mixture of Experts (MoE) architectures. While MoE improve parameter efficiency, effectively applying MoE to simultaneously model modality-specific features and cross-modal associations in LVLMs remains challenging. In this work, we propose to incorporate Mixture of Intra- and Inter-Modality Experts (MoIIE) to LVLMs. For each token, expert routing is guided by its modality, directing tokens to their respective intra-modality experts as well as a shared pool of inter-modality experts, enabling the model to jointly learn rich intra-modal features and cross-modal interactions. We further introduce an effective and straightforward two-stage training strategy, which facilitates the direct activation of both MoE and multi-modal capabilities. Extensive experiments across different data scales, visual encoder and LLM backbone demonstrate the effectiveness, efficiency and generality of our approach. Notably, our MoIIE models with 5.5B and 11.3B activated parameters match or even surpass the performance of existing advanced open-source MoE-LLMs based multi-modal models that involve more activated parameters.
MoIIE: Mixture of Intra- and Inter-Modality Experts for Large Vision Language Models
Dianyi Wang1,2  Siyuan Wang3  Zejun Li1  Yikun Wang1,2 Yitong Li4  Duyu Tang4  Xiaoyu Shen5  Xuanjing Huang1  Zhongyu Wei1,2 1Fudan University  2Shanghai Innovation Institut 3University of Southern California  4Huawei Technologies Co., Ltd 5Ningbo Key Laboratory of Spatial Intelligence and Digital Derivative, Institute of Digital Twin, EIT dywang24@m.fudan.edu.cn, sw_641@usc.edu
@@ 1 Introduction
To this end, we proposes the Mixture of Intra- and Inter-Modality Experts (MoIIE), to simultaneously capture both modality-specific features and cross-modal associations in LVLMs. As illustrated in Figure 1 (b), MoIIE comprises three distinct groups of experts: two intra-modality groups that specialize in language and vision, independently processing text and image tokens; and an inter-modality expert group shared by both modalities that focuses on cross-modal interactions between text and image tokens. Correspondingly, MoIIE learns two dedicated routers respectively for image and text tokens, with each token dynamically routed to the most relevant experts from the corresponding intra-modality group and the inter-modality group. As not all tokens are strongly associated across modalities, such as function words in text or visual regions without descriptive content, the routing mechanism allows for flexible combinations. Some tokens may activate only intra-modality experts or only inter-modality experts, while others may access both expert groups simultaneously. This dynamic routing strategy is consistently applied to both image and text tokens. Additionally, for multi-modal MoE training, most existing approaches adopt a sparse upcycling strategy Komatsuzaki et al. (2023) that transforms dense LVLMs into sparse models. This strategy generally follows a three-stage training pipeline: multi-modal pre-training of the dense model, fine-tuning the dense model on diverse multitasking datasets with all parameters unfrozen, and model sparsification with only expert modules trained Lin et al. (2024a); Chen et al. (2024b). However, this pipeline is relatively cumbersome and the limited training during the final stage potentially undermines the model’s generalization capabilities. In contrast, we propose that multi-modal fine-tuning and sparsification can be jointly optimized with all parameters updated in a more straightforward and effective two-stage training strategy. In the first stage, visual inputs are aligned with the LLM backbone through multi-modal pre-training datasets. In the second stage, MoIIE is integrated and the entire model is optimized using fine-tuning datasets, to simultaneously learn (1) versatile multi-modal understanding and instruction-following capabilities, while enabling (2) expert learning of modality-specific knowledge with cross-modal association within the MoIIE module. Our experimental results demonstrate that this simplified strategy is easy to implement and also effectively enhances performance across diverse downstream tasks.
@@ 2.2.2 Sparse MoIIE Forward
$\text{MoIIE}(x_{n})=\begin{cases}\sum_{i=1}^{K}G_{i}^{\mathcal{T}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{T}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{T}\\ \sum_{i=1}^{K}G_{i}^{\mathcal{I}}(x_{n})\cdot\text{E}_{i}^{\ \mathcal{I}/S}(x_{n}),&\text{if }x_{n}\in\mathcal{I}\end{cases}$
where each token $x_{n}$ is processed by its top-$K$ activated experts $\text{E}_{i}(x_{n})$, with their outputs aggregated through a probability-weighted summation. Specifically, image tokens are routed by the visual router $G^{\mathcal{I}}(x)$ to the top-$K$ experts selected from the combined pool of intra-vision and inter-modality groups $E^{\mathcal{I/S}}$, while textual tokens are routed by the textual router $G^{\mathcal{T}}(x)$ to the top-$K$ experts from the intra-language and inter-modality groups $E^{\mathcal{T/S}}$. The routing mechanism enables flexible expert selection patterns: each token may activate only intra-modality experts, only inter-modality experts, or a combination of both. This hierarchical routing strategy ensures efficient multi-modal learning while simultaneously modeling both modality-specific features and cross-modal associations. All experts in $\mathcal{E}$ are initialized from the LLM’s FFN layers, preserving prior knowledge while enabling adaptive multi-modal specialization.
@@ 3.1 Implementation Details
We employ a pre-trained SigLIP model Zhai et al. (2023b) and CLIP Radford et al. (2021)in Appendix A.4 as visual encoder, a two-layer MLP as connection module. For the LLM backbone, we utilize phi-3-mini Marah Abdin (2024) and LLaMA3-8b AI@Meta (2024). Our method and all compared MoE architectures use a 4-experts configuration. Specifically, our MoIIE module includes 2 intra-modality experts (one for vision and one for language) and 2 inter-modality experts. More detailed discussion of expert settings is provided in the ablation studies 3.6.
@@
[…]
@@
[…]
During training, we utilize the AdamW optimizer Kingma and Ba (2017) with a cosine learning rate scheduler for one epoch across both stages. In the first stage, only the connection module is optimized with a learning rate of $1\times 10^{-3}$, using the Bunny-pretrain-LAION-2M dataset He et al. (2024). In the second stage, all parameters are unfrozen for joint SFT and sparse upcycling training. The learning rate is set to $2\times 10^{-6}$ for the visual encoder and $2\times 10^{-5}$ for all other components. For this stage, we attempt training with different scales of visual instruction data: (1) 1.3M samples from MGM-Instruction Li et al. (2024c); (2) 2M samples by adding Bunny-695K He et al. (2024); (3) 2.7M samples by further incorporating LLaVA-NEXT-779k Liu et al. (2024b)(4)8M advanced training data from LLaVA-OV Li et al. (2024a). Training is conducted using DeepSpeed Shuaiwen Leon Song (2023) with ZeRO-3 optimization. All compared baselines in the main results and our method follow the same training setup and data configurations. Further implementation details and results are provided in Appendix A.
@@ 3.2 Evaluation Benchmarks
We conduct a comprehensive evaluation across 13 diverse multi-modal benchmarks. Specifically, general multi-modal benchmarks include MMBench-EN Liu et al. (2024c), MM-Vet Yu et al. (2024), GQA Hudson and Manning (2019), VQAv2 and SEED-Image  Li et al. (2023a). For knowledge-based question answering, we utilize MMMU validation subset Yue et al. (2024), AI2D Kembhavi et al. (2016),SciQA-IMG Lu et al. (2022) and Mathvista Lu et al. (2024b). For OCR-based question answering, we assess performance on TextVQA Singh et al. (2019), ChartQA Masry et al. (2022), and DocVQA Mathew et al. (2021). Additionally, we evaluate hallucination robustness using POPE Li et al. (2023c) and HallusionBench Guan et al. (2024).
@@ 3.3 Main Results
We compare our MoIIE-equipped model against three baseline variants: the original dense model (Dense) without MoE, the vanilla MoE model MistralAITeam (2023) that routes tokens from all modalities to a shared pool of experts (Vanilla MoE), and the model with modality-specific experts where text and image tokens are routed to their respective specialized experts (Modality MoE). The comparisons are conducted across multiple visual instruction tuning data scales and varying LLM backbones, as shown in Table 3.1. The key findings are as follows:
(1) The MoIIE module consistently outperforms all other architectural variants across most benchmarks with different data scales, visual encoders, and LLM backbone sizes, demonstrating the powerful generalization capability of our MoIIE. Integrating both modality-specific features and cross-modal associations can effectively enhance performance on diverse multi-modal tasks. Further analysis of expert modules is provided in Appendix A.3.
@@
(3) We further provide detailed performance trends with increasing fine-tuning data in Figure 3. As the data scale continues to increase, our MoIIE progressively improves while other architectures, especially Dense and Modality MoE, encounter performance limitations. Moreover, with larger training datasets, our MoIIE consistently outperforms all alternatives, with the performance gap between MoIIE and other architectures widening. This suggests that the MoIIE framework offers superior scaling properties for multi-modal learning, effectively leveraging larger datasets to enhance representation power without the parameter inefficiency of dense models or the limited cross-modal reasoning of strictly modality-separated experts.
@@ 3.4 Comparison with MoE-based LLMs Initialized LVLMs
We compare our MoIIE, pre-trained on the LLaVA-OV-Single-Image dataset Li et al. (2024a), with open-source LVLMs initialized from MoE-based LLMs. Models such as SPHINX-X Liu et al. (2024a), MGM Li et al. (2024c), and CuMO Li et al. (2024b), built on Mixtral-8x7B MistralAITeam (2023), exploit large-scale language data for robust initialization but require substantial computational resources, limiting practicality in constrained settings. By contrast, MoIIE attains comparable performance across diverse multimodal benchmarks with 17% fewer activated parameters, thereby reducing both training and deployment costs. Moreover, it enables constructing MoE-based LVLMs from arbitrary LLM backbones, offering flexible expert configurations without performance degradation.
@@
[…]
@@
[…]
@@ 3.5 The Discussion of Training Pipeline
Converting dense LLMs into MoE-based LVLMs typically follows a three-stage pipeline Lin et al. (2024a); Chen et al. (2024b): (1) multimodal pretraining, (2) supervised fine-tuning (SFT) with multimodal instructions, and (3) sparse upcycling to obtain MoE-based LVLMs. This process is complex, and limited training in the final stage often weakens generalization. We propose a simplified two-stage approach: after pretraining, SFT and sparsification are combined into one phase, enabling direct activation of both multimodal and MoE features. Specifically, the three-stage pipeline uses 2M visual instruction samples for SFT (tuning all parameters) and LLaVA-NEXT-779k for sparse upcycling (only tuning MoE layers), while our two-stage method uses 2.7M samples for joint SFT and sparse upcycling (tuning all parameters). As shown in Table 3, our approach consistently outperforms the three-stage pipeline under the same conditions, offering both simplicity and improved effectiveness.
@@ 3.6 Ablation Study
Increasing the number of experts generally enhances MoE performance Lepikhin et al. (2020); FedusF et al. (2022). As shown in Table 4, expanding the MoIIE expert pool from 4 to 8 yields gains of 0.3, 0.2, and 0.9 points on MMBench, GQA, and MMMU, respectively, while results on POPE and TextVQA remain stable or slightly decline. These findings suggest that larger expert pools improve generalization on complex benchmarks but may require more data to fully realize their capacity.
@@
Maintaining a balanced allocation of intra- and inter-modality experts is crucial for performance. As shown in Table 4, under 8-expert setup, Balanced configuration (2 vision, 2 language, 4 shared; 1:1 ratio) consistently outperforms the Unbalanced configuration (3 vision, 3 language, 2 shared; 3:1 ratio). The balanced design ensures more effective cross-modal interaction, getting stronger overall performance.
@@
We investigate two configurations for integrating MoIIE into LVLMs: an interleaved design, where MoIIE modules are inserted every other layer alongside dense layers, and a full design, where all layers are replaced with MoIIE modules Lin et al. (2024a); AI@Meta (2025). As shown in Table 4, the full design does not yield notable performance gains but significantly increases training cost and computational overhead due to the larger parameter count. By contrast, the interleaved configuration achieves comparable results with substantially higher efficiency and lower resource usage. We therefore adopt the interleaved design as the optimal balance of performance, efficiency, and scalability for MoIIE-based LVLMs.
Impact of Load Balance Coefficient
In MoE architectures, there is a trade-off between load balancing and performance. While the auxiliary loss promotes balanced expert usage Cai et al. (2024), excessive balancing can hinder multimodal learning and weaken cross-modal associations. As shown in Table 4, increasing the auxiliary loss weight $\alpha$ consistently reduces performance, especially on MMMUval and MM-Vet. These results underscore the importance of tuning $\alpha$ to maintain expert specialization and robust multimodal understanding.